Group Selection Page HELP Controls Group ITDM Vacuum WP78 SPB SQS XO ToBeMigrated2XWiki ToBeMigrated2GitlabPages ToBeMigratedByGroup Migrated2Zulip Archived2PDF not2MigrateOrArchive
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID PSPO_XO XO_Operation SQS Migrated2Zulip FXE Migrated2Zulip DET Sample Environment SPB Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2023 logbook, Contact: F.Wolff-Fabris, A.Galler, L.Samoylova, H.Sinn  Not logged in ELOG logo
Message ID: 85     Entry time: 17 Mar 2023, 01:48
Author: Johannes Moeller 
Group: MID 
Subject: Shift summary 

The intensity of the Seeding decreesed from 850 uJ to 550 uJ during the time between 20:00 and 02:00.

The Spectral purity also got worst, the beam in HIREX looks to have more SASE shoulders.

Between 18 and 20 we show a prossible transiant signal, PPL, of the lower branch. But a problem with the monochromator in the MID opt Hutch trigger the valve to close. After we where not able to see the signal possible effect of laser or X-ray inestabilities.

Problems with the liquidjet, the nozzel got clogged.

Started the aligment of the upper branch of SDL

Attachment 1: Screenshot_from_2023-03-17_01-44-53.png  77 kB  | Hide | Hide all
Screenshot_from_2023-03-17_01-44-53.png
ELOG V3.1.4-7c3fd00